OpenStudy (anonymous):

How would you integrate the following function? Particularly, what would you do after the last following step? integral( (9-x^2)^(1/2) dx) x = 3 sin u u = invsin(x/3) integral( 3 (cos u) d?

OpenStudy (anonymous):

fuk

OpenStudy (anonymous):

I have the answer sheet, and wolfram gives a decent explanation, but I can't understand what the next step is.

OpenStudy (anonymous):

pull out 3 get 3*integral(cos u du) = 3sin u + C

OpenStudy (anonymous):

You just have to know the integral of cos(u)

OpenStudy (anonymous):

Nah, it's still in the form dx, not du. My worksheet says the answer is. (1/2)x (9-x^2)^(1/2) + 9 invsin(x/3))

OpenStudy (anonymous):

Once you integrate it you can plug back in what you substituted for u

OpenStudy (anonymous):

You need to solve for dx in terms of du.

OpenStudy (anonymous):

plug u back in

OpenStudy (anonymous):

Yeah, that's what I'm having issues with. I've never been strong with trig functions.

OpenStudy (anonymous):

When you plug u in, there's a sin and cosine function, which indicates integration by parts, that seem right?

OpenStudy (anonymous):

\[x = 3 sin(u) \implies dx = 3cos(u)du\] \[(9-x^2)^{1/2} dx= [9-9cos^2(u)]^{1/2} \cdot (3cos(u)\ du)\]

OpenStudy (anonymous):

Then you can factor out a 3 from the sqrted factor and a 3 from the du part and have a 9 out front and rewrite it as \((1-cos^2(u))^{1/2}cos(u) du\). This solves nicely with another substitution.

OpenStudy (anonymous):

oh wait, no.

OpenStudy (anonymous):

It's a tough one XD

OpenStudy (anonymous):

u use sin(x)^2 + cos(x)^2 = 1

OpenStudy (anonymous):

yeah, 1-cos^2(u) = sin^2(u)

OpenStudy (anonymous):

which you can take the square root of easily and also u-sub easily.

OpenStudy (anonymous):

so z = sin u dz = cos x dx

OpenStudy (anonymous):

i mean dz = cos u du

OpenStudy (anonymous):

integral(z) = z^2/2 then replace with u and then replace with x

OpenStudy (anonymous):

I have ... \[3\cos(\sin^-(x/3))\] Where would I go from here?

OpenStudy (anonymous):

that's not right. How did you get there?

OpenStudy (anonymous):

\[\cos x =(1- \sin^2 x)^{?}\]

OpenStudy (anonymous):

other way round bob.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x = 3sin(u) x/3 = sin(u) invsin(x/3) = u integral ( 3cos(u)) integral (3cos(invsin(x/3)))

OpenStudy (anonymous):

you don't want to sub back in your u until after you integrate.

OpenStudy (anonymous):

so that'll be negative sin(u)

OpenStudy (anonymous):

because the integral of cosine is sin, right? See, I'm confused because it's still technically in the form dx, not du.

OpenStudy (anonymous):

negative sin*

OpenStudy (anonymous):

It shouldn't be. When you do the u-sub you need to replace your dx with some function of u du

OpenStudy (anonymous):

This is what you're starting with right? \[\int (9-x^2)^{1/2} dx\]

OpenStudy (anonymous):

Yessir.

OpenStudy (anonymous):

Ok, and we let \(x = 3sin(u)\).

OpenStudy (anonymous):

So you know, I'm studying for a test tomorrow, this isn't homework.

OpenStudy (anonymous):

yeah, np

OpenStudy (anonymous):

Ok, so now what is dx in terms of du?

OpenStudy (anonymous):

dx = 3cosu du

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

So now lets swap out \(x^2\) and \(dx\) with the appropriate expressions of u and du.

OpenStudy (anonymous):

What do you have?

OpenStudy (anonymous):

(9-9sin^2u)^(1/2) dx (9(1-sin^2u))^(1/2) dx 3 cos u dx

OpenStudy (anonymous):

you didn't swap out dx.

OpenStudy (anonymous):

I don't know how the hell to do that my friend :P

OpenStudy (anonymous):

dx = 3cosu du You said it yourself.

OpenStudy (anonymous):

Agreed, but if you .... OH!

OpenStudy (anonymous):

3cosu * 3 cosu ?

OpenStudy (anonymous):

Indeed (along with a du)

OpenStudy (anonymous):

okay, so we have ... \[\int\limits_{}^{} 3 \cos (u)cos(u) du\] which is ... \[\int\limits_{}^{} 9 \cos ^2 (u) ^ {1/2} du\]

OpenStudy (anonymous):

forget that square root

OpenStudy (anonymous):

no its Integral (3 cos^2 x)

OpenStudy (anonymous):

Right. So then you can use the double/half angle equation to swap the \(cos^2\) for an expression involving the cos(2u)

OpenStudy (anonymous):

She said she would be providing those equations on the test.

OpenStudy (anonymous):

Should be 9 cause there's a 3 from the 9 under the sqrt and a 3 from the du replacement of dx.

OpenStudy (anonymous):

\[\cos^2 x = (1 + \cos 2x )/2\]

OpenStudy (anonymous):

that's the one.

OpenStudy (anonymous):

I think polpak's solution is about the same as the attachment.

OpenStudy (anonymous):

Yeah but polpak is a better teacher than any pdf :p

OpenStudy (anonymous):

I'm working out the cos squared substitution on paper as we speak.

OpenStudy (anonymous):

particularly a pdf that is a screenshot of wolframalpha.

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

Though alpha is good for checking your answers and finding where you made a mistake, but sometimes it's process is a little cumbersome.

OpenStudy (anonymous):

its rather. I hate that stupid part of english.

OpenStudy (anonymous):

Tell me about it, I checked wolfram and couldn't find how they went from 3cos(u) to 9cos^2(u). This is getting uglier and uglier as we go along. I have... \[\int\limits_{}^{} 9 {( 1+\cos2u)}/2 du\]

OpenStudy (anonymous):

That's not ugly at all!

OpenStudy (anonymous):

Break it into two integrals.

OpenStudy (anonymous):

and pull the 9/2 all the way out front.

OpenStudy (anonymous):

Oh wow. You're right. Hah! I need to practice these more, for sure.

OpenStudy (anonymous):

Then the only thing you have to do is another sub for z = 2u on the second integral if you don't feel confident doing it in your head.

OpenStudy (anonymous):

Nah it's just division by 2

OpenStudy (anonymous):

indeed.

OpenStudy (anonymous):

\[9/2 ( u - \sin(2u)/u)\] Does this look alright?

OpenStudy (anonymous):

why is that u under the sin(2u) ?

OpenStudy (anonymous):

Oh, sorry that's supposed to be a 2. I have dyslexia, it's a pain in the rear end sometimes.

OpenStudy (anonymous):

yes. Me too.

OpenStudy (anonymous):

With a 2 in the denominator that is correct.

OpenStudy (anonymous):

oh, wait

OpenStudy (anonymous):

shouldn't be a -sin(2u)

OpenStudy (anonymous):

\[\int cos(u)du = sin(u) +C \]

OpenStudy (anonymous):

should be positive.

OpenStudy (anonymous):

the derivative of cos is -sin, but the integral of cos is just sin

OpenStudy (anonymous):

\[u = \sin^{-1}(x/3)\] You're right, I got that mixed up with the rules for sin integration.

OpenStudy (anonymous):

yeah, so that works nicely given what u is.

OpenStudy (anonymous):

So if I were to substitute u, then I would be finished?

OpenStudy (anonymous):

you can simplify probably, but it's up to you

OpenStudy (anonymous):

One quick note, in my packet, the answer is 1/2 (x * (9-x^2)^(1/2) + 9 sin -1 (x/3)) How can it be so different?

OpenStudy (anonymous):

because they plug back in for sin(2u) and use the formula for \(sin(2\theta)\)

OpenStudy (anonymous):

to just get a sin or cos squared which they then swap out for 1-cos or sin squared

OpenStudy (anonymous):

all kinds of fun trig subs going on.

OpenStudy (anonymous):

We can work through it if you want

OpenStudy (anonymous):

well, I'm happy leaving it in the form inverse sine. I'm just curious what would happen when you substitute sin^-1(x/3) into sin(2u)/2

OpenStudy (anonymous):

because I believe she would want it back in the form of x

OpenStudy (anonymous):

I would imagine black holes are created, or something of that variety (when sin and inverse sin collide)

OpenStudy (anonymous):

No, you just get back what you started with. \[sin(sin^{-1}(x/3)) = x/3\]

OpenStudy (anonymous):

What happened to the two?

OpenStudy (anonymous):

it's still there, I'm showing a different example

OpenStudy (anonymous):

Ok, so here it is. \[\frac{9}{2}sin(2u) =\frac{ 9(2)}{2}cos(u)sin(u)\]\[ =9(\sqrt{1-sin^2(u)})sin(u) \]